Zeronknight (zeronknight):

Match the inequality with its solution. -\frac{2}{5}x> 14 A. x > 35 B. x < 35 C. x > -35 D. x < -35

Zeronknight (zeronknight):

\[-\frac{2}{5}x> 14\]

pooja195 (pooja195):

Mutliply by the reciporcal

pooja195 (pooja195):

\[(-5/2)-\frac{2}{5}x> 14(-5/2)\]

OpenStudy (mathmale):

I agree with Pooja. The reciprocal of -2/5 is -5/2. Note that if you multiply an inquality by a negative quantity, you MUST reverse the direction of the inequality symbol.

pooja195 (pooja195):

@mathmale my plan was to do that at the end because I did not want to complicate things.

Zeronknight (zeronknight):

owo????

pooja195 (pooja195):

\[(-5/2)-\frac{2}{5}x> 14(-5/2)\] What is 14(-5/2)=?

Zeronknight (zeronknight):

Not so sure?

pooja195 (pooja195):

\[\huge~\rm~\bf~\frac{ 14\times -5=? }{ 1\times 2=? }\]

Zeronknight (zeronknight):

should 14 be a negative? and 1 x 2 = 2

Zeronknight (zeronknight):

Well would 14 x -5 = -insert negative number here?-

pooja195 (pooja195):

Do 14 x 5 then add the negative sign

Zeronknight (zeronknight):

-70 then

pooja195 (pooja195):

Good -70 divided by 2=?

Zeronknight (zeronknight):

35

OpenStudy (anonymous):

- 35

pooja195 (pooja195):

when we divide a positive by a negative the result will be a negative

pooja195 (pooja195):

So... x>-35 what do we get once we flip the sign?

Zeronknight (zeronknight):

X < -35

pooja195 (pooja195):

Good job well done :-)

Zeronknight (zeronknight):

But 70 was positive?

Zeronknight (zeronknight):

Oh nevermind XD

pooja195 (pooja195):

I mentioned above when you divide a postive by a negative vice versa you get a negative answer

Zeronknight (zeronknight):

When we times a postivie and a negative the result is negative. When we divide a positive by a negativee we get negative? @pooja195